Circular No. 8620

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2005 RV_{25} (LONEOS-CHRISTENSEN) A comet was reported by E. J. Christensen (Catalina Sky Survey, 0.68-

m Schmidt telescope) as showing a 15′′ coma and no discernible tail in four stacked 60-second exposures on Oct. 22.3 UT. It was confirmed as cometary on CCD images taken by R. H. McNaught (Siding Spring Observatory, 1.0-m f/8 reflector; fan tail spanning 90^{o} to the southwest) on Oct. 22.6, by P. Birtwhistle (Great Shefford, Berkshire, England, 0.40-m f/6.0 Schmidt-Cassegrain reflector; 10′′ coma, moderately condensed, no tail) on Oct. 22.9, by R. Miles (Stourton Caundle, Dorset, England, 0.28-m f/10 Schmidt-Cassegrain reflector) on Oct. 23.0, by J. Young (Table Mountain Obsrvatory, 0.6-m f/16 Cassegrain reflector; very diffuse 8′′-diameter coma with little central condensation and a very broad but poorly defined tail spanning p.a. 170^{o}–250^{o} extending to ∼ 20′′ from the coma center) on Oct. 23.2, and by J. E. McGaha (Tucson, AZ, 0.62-m f/5.1 reflector; elongated coma of size 8′′ × 12′′ toward p.a. 225^{o}) on Oct. 23.2. Computations at the Minor Planet Center revealed several prediscovery observations, notably that the object is identical with the two-night asteroidal object 2005 RV_{25} (LONEOS discovery, MPS 141815). Discovery positions follow, together with the orbital elements given on MPEC 2005-U18, which also contains further astrometry and an ephemeris.

2005 UT α_{2000} δ_{2000} Mag. Observer
Sept.11.28012 1^{h}22^{m}06.^{s}33 +19^{o}05′16′′.3 17.7 LONEOS
Oct. 22.25151 1 00 20.98 +16 33 08.1 18.3 Christensen
T = 2006 Nov. 11.2775 TT  ω = 192.^{o}3492
e = 0.167169  Ω = 246.8919 }2000.0
q = 3.604711 AU  i = 9.8917
a = 4.328260 AU n^{o} = 0.1094546 P = 9.005 years

2005 SD Further to IAUC 8618, several CCD observers have reported that their

images of 2005 SD show no cometary appearance at all, including C. Hergenrother (Mt. Graham 1.8-m VATT Lennon telescope, co-added 6-min exposure on Oct. 3.28 UT, and Mt. Hopkins 1.2-m reflector, co-added 300s R-band image on Oct. 22.07), R. Ferrando (Valencia, Spain, 0.36-m f/6.3 Schmidt-Cassegrain reflector, Oct. 21.86 and 23.78), and J. Young and D. Mayes (Table Mountain, 8-min exposures on Oct. 23.2).

2005 October 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green